yourgrandpa:

in 2000, the population of Seattle was 3,043,878. by 2015, the population increased to 3,733,580. what was the percent increase in the population of Seattle between 2000 and 2015 rounded to the nearest hundredth? A)18.47% C)80.53% D)122.66%

AZ:

We answered this question yesterday. Was it wrong too??

yourgrandpa:

yes and actually the other question was right my bad the teacher realized it was correct sorry for the inconvenience

AZ:

Maybe the teacher will realize she made a mistake with this one too :P \(\text{% change} = \dfrac{\text{new value - original}}{\text{original}} \times 100\text{%}\)

yourgrandpa:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @AZ Maybe the teacher will realize she made a mistake with this one too :P \(\text{% change} = \dfrac{\text{new value - original}}{\text{original}} \times 100\text{%}\) \(\color{#0cbb34}{\text{End of Quote}}\) lol probably

jhonyy9:

so like a first step subtract from 3733580 the 3043878 suppose you get P after this just use the rule of 3 simple and write in this way 3043878 mean 100% P mean x% cross multiplie and calcule the x value x = P*100/3043878 = ?%

AZ:

What is answer choice B? I only see A, C, and D

yourgrandpa:

B was wrong

AZ:

So it was the 22.6%

yourgrandpa:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @jhonyy9 so like a first step subtract from 3733580 the 3043878 suppose you get P after this just use the rule of 3 simple and write in this way 3043878 mean 100% P mean x% cross multiplie and calcule the x value x = P*100/3043878 = ?% \(\color{#0cbb34}{\text{End of Quote}}\) 689702

AZ:

Well I guess like that previous question, we'll just have to divide by the other number (even though your teacher realized she made a mistake) So good, you have 689,702 what is 689,702 / 3,733,580

jhonyy9:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @yourgrandpa \(\color{#0cbb34}{\text{Originally Posted by}}\) @jhonyy9 so like a first step subtract from 3733580 the 3043878 suppose you get P after this just use the rule of 3 simple and write in this way 3043878 mean 100% P mean x% cross multiplie and calcule the x value x = P*100/3043878 = ?% \(\color{#0cbb34}{\text{End of Quote}}\) 689702 \(\color{#0cbb34}{\text{End of Quote}}\) this is P ? then yes so use what i ve wrote above for the next step

yourgrandpa:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @AZ Well I guess like that previous question, we'll just have to divide by the other number (even though your teacher realized she made a mistake) So good, you have 689,702 what is 689,702 / 3,733,580 \(\color{#0cbb34}{\text{End of Quote}}\) 0.184729

AZ:

And don't forget to multiply by 100 to get it as a percent

yourgrandpa:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @AZ And don't forget to multiply by 100 to get it as a percent \(\color{#0cbb34}{\text{End of Quote}}\) 18.472900

jhonyy9:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @jhonyy9 \(\color{#0cbb34}{\text{Originally Posted by}}\) @yourgrandpa \(\color{#0cbb34}{\text{Originally Posted by}}\) @jhonyy9 so like a first step subtract from 3733580 the 3043878 suppose you get P after this just use the rule of 3 simple and write in this way 3043878 mean 100% P mean x% cross multiplie and calcule the x value x = P*100/3043878 = ?% \(\color{#0cbb34}{\text{End of Quote}}\) 689702 \(\color{#0cbb34}{\text{End of Quote}}\) this is P ? then yes so use what i ve wrote above for the next step \(\color{#0cbb34}{\text{End of Quote}}\) substitute there in place of P

AZ:

\(\color{#0cbb34}{\text{End of Quote}}\) this is P ? then yes so use what i ve wrote above for the next step \(\color{#0cbb34}{\text{End of Quote}}\) We had tried it that way with 3,043,878 but we got it wrong https://questioncove.com/users/yourgrandpa#/updates/6033f4b14ec784025b12d80

jhonyy9:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @AZ And don't forget to multiply by 100 to get it as a percent \(\color{#0cbb34}{\text{End of Quote}}\) exactly

AZ:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @yourgrandpa 18.472900 \(\color{#0cbb34}{\text{End of Quote}}\) So A!

yourgrandpa:

thank youuuuuu so much yourr a big help to me (:

AZ:

Anytime!